Sct up -accounlts lor each account and post the closing entries. 3-42. Analyzing and Reporting Financial Statement Effects of Transactions M. E. Carter launched Carter Company, a professional services firm on March 1. The firm will prepare financial statements at each month-end. In March Gits first month), Carter executed the following transactions. Enter the transactions, a through g, into the financial statement effects template shown in the module. a. Carter (owner) invested in the company $100,000 cash and $20,000 in property and equipment. The company issued common stock to Carter. b. The company paid $3,200 cash for rent of office furnishings and facilities for March. c. The company performed services for clients and immediately received $4,000 cash earned. d. The company performed services for clients and sent a bill for $24,000 with payment due within 60 days. e. The company compensated an office employee with $4,800 cash as salary for March. f. The company received $10,000 cash as partial payment on the amount owed from clients in transaction d. &. The company paid $935 cash in dividends to Carter (owner).
Answer #1
Transaction Cash asset + Non cash asset = Liabilities + Contrib. Capital + Earned capital
a. Issued stock for cash and PPE 100,000 + 20,000 = 0 + 120,000.00 + 0
b. Paid rent -3,200 + 0 = 0 + 0 + -3,200
c. Performed services for cash 4,000 + 0 = 0 + 0 + 4,000
d. Performed services on account 0 + 24,000 = 0 + 0 + 24,000
e. Paid wages -4,800 + 0 = 0 + 0 + -4,800
f. Received payments on account receivable 10,000 + -10,000 = 0 + 0 + 0
g. Paid dividends -935 + 0 = 0 + 0 + -935